OpenStudy (anonymous):

20. 5d = 17.1 20. 5d = 17.1 @Mathematics

OpenStudy (chriss):

\[\frac{5d}{5} = \frac{17.1}{5}\ d = 3.42

OpenStudy (amistre64):

lost a ] on the end ;)

OpenStudy (chriss):

\[\frac{5d}{5} = \frac{17.1}{5}\]\ d = 3.42